TASK SECONDS TASKS THAT MUST PRECEDE A 25 — B 15 A C 20 B D...

TASK

SECONDS

TASKS THAT
MUST PRECEDE

A

25

B

15

A

C

20

B

D

29

B

E

16

C

F

8

D

G

35

E, F

H

32

G


The workday is 10.5 hours long. Demand for completed product is 900 per day.

a. Find the cycle time required to meet the desired output rate. (Round your answer to 1 decimal place.)

b. What is the theoretical number of workstations required to meet the desired output rate? (Round up your answer to the next whole number.)

d. Balance the line using sequential restrictions and the longest-operating-time rule. (Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 1 decimal place.)

e. What is the efficiency of the line balanced as in part (d) assuming it is running at the cycle time determined in part (a)? (Round your answer to 1 decimal place.)

f. Suppose that demand rose from 900 to 950 units per day. What would you do? (Round your answer for cycle time down to the nearest whole number.)

g. Suppose that demand rose from 900 to 1,150 units per day. What would you do? (Round your answer for cycle time down to the nearest whole number.)
Answer #1

Answer a: desired output = 900 units per 10.5 hours= 10.5*60*60/240= 42 secs per unit

Hence the workstation cycle time required is 42 secs per unit

Total task time =25+15+20+29+16+18+35+32= 190 secs

Minimum no. of work station = total task time/ cycle time = 190/42 =4.5 ~5

Answer b: Theoretical minimum no. of workstation is 5

Answer c: Line balancing, required cycle time is 42 secs per workstation. Sum of task times in a workstation cannot be more than cycle time.

Workstation

Tasks

Idle time= Cycle time- sum of task time

Reason

1

A, B

42-25-15= 2

Based on precedence

2

D, F

42-29-8= 5

D is the longer task than C

3

C, E

42-16-20= 6

There was no space for C after D

4

G

42-35= 7

5

H

42-32= 10

Answer d: Total idle time= 2+5+6+7+10= 30 secs

Balanced delay = total idle time/ no. of workstation* cycle time = 30/(5*32)= 0.1923 ie 19.23%

Efficiency= 1-balanced delay= 80.77%

Note: Chegg's guidelines mandate to solve one question or 1st four subparts of a question at a time. Please post again for remaining questions.